Given the schematic notation: Cu | Cu2+ (0.0200 M) ∥ Ag+


(0.0200 M) | Ag, calculate the cell


potential of this system. Show all calculation.

1
Expert's answer
2021-11-18T10:40:01-0500

E cell "=" E right "-" E left


Ag+ + e- "\to"Ag(s) E"\theta=0.799v"


Cu2+ +2e- "\to" Cu(s) E"\theta=0.337v"



E"_{Ag^+}/_{Ag}=0.799-0.0592 \>log \>\>\frac{1}{0.0200}=0.6984v"



E "_{Cu^2+}/_{Cu}=0.337-\frac{0.0592}{2}\>log\frac{1}{0.0200}=0.2867v"



E "_{cell}=" E"_{Ag^+}-" E"_{Cu^{2+}}=0.6984-0.2867"


"=0.412v"